OpenStudy (anonymous):

PLEASE HELP! Which of the following functions have a domain of (-∞,∞) and a range of (5,∞)? Check all that apply. a. f(x)=3^x+5 b. f(x)=3^x-5 c. f(x)=0.25^x+5 d. f(x)=0.25^x-5

OpenStudy (lucaz):

all of them has the same domain (-∞,∞), but the in first one, if you replace x by positive infinity you will get positive infinity, and replacing by negative intinity you wil get 1/(3^∞) + 5, wich is something really really small plus 5, this is the range (5,∞)

OpenStudy (anonymous):

so the answers are b,c,and d?

OpenStudy (lucaz):

it's A, did you understand what I said before?

OpenStudy (anonymous):

yeah it was just confusing which one you were talking about

OpenStudy (lucaz):

right.. it's the first one

OpenStudy (anonymous):

are you sure the others don't work because I can have multiple answers

OpenStudy (lucaz):

let's see if they work, the 2nd option 3^x-5 3^∞-5 = ∞ 3^-∞ -5 = 1/(3^∞) -5 something really smal, almost zero, minus five, this is almost equal to negative five -5, the range would be (-5,∞)

OpenStudy (anonymous):

so no

OpenStudy (lucaz):

the other two options are kind of the inverse

OpenStudy (lucaz):

0.25 is the same as 1/4 right?

OpenStudy (anonymous):

yes

OpenStudy (lucaz):

(1/4)^∞ = 1/(4^∞) = almost zero (1/4)^-∞ = 4^∞/1 = infinity

OpenStudy (lucaz):

the correct options are A and C

OpenStudy (anonymous):

thank you

OpenStudy (lucaz):

you're welcome